OpenStudy (starryknight):

What is the range of the following function? f(x) = √x - 3 + 2 x ≥ 2 f(x) ≥ 2 x ≥ 3 f(x) ≥ 3 please help me i dont understand this

OpenStudy (anonymous):

which two did you narrow it down to?

OpenStudy (starryknight):

the 2 on the right side

OpenStudy (anonymous):

ok good

OpenStudy (starryknight):

?

OpenStudy (anonymous):

now it is very similar to the last one in this case \[\sqrt{whatever}\geq 0\] always so \[\sqrt{whatever}+2\geq 2\] that is all